OpenStudy (anonymous):

What is the equation of a parabola that has a vertical axis, passes through the point (–1, 3), and has its vertex at (3, 2)? x= –y2/16+6y/16–41/16 y= –x2/16+6x/16–41/16 y=x2/16–6x/16+41/16 x=y^2/16–6y/16+41/16

OpenStudy (anonymous):

x= –y^2/16+6y/16–41/16 y= –x^2/16+6x/16–41/16 y=x^2/16–6x/16+41/16 x=y^2/16–6y/16+41/16

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

\[x=-\frac{ x^2 }{ 16 }+\frac{ 6y }{ 16}-\frac{ 41 }{ 16}\]

OpenStudy (anonymous):

Thats how a is set up

OpenStudy (isaiah.feynman):

Okay.

OpenStudy (isaiah.feynman):

The standard equation of a parabola is given by

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

getting sleepy

OpenStudy (anonymous):

trying to fight it

OpenStudy (anonymous):

been waiting all night

OpenStudy (isaiah.feynman):

Waiting for what? lol

OpenStudy (anonymous):

I'm going to find this answer before I go to sleep and for the day over with

OpenStudy (michele_laino):

if I write your parabola as below: y=ax^2+bx+c then I can write this: -b/2a=-1

OpenStudy (anonymous):

so It is neither A or D

OpenStudy (michele_laino):

sorry, can you check your data?

OpenStudy (michele_laino):

if the vertical axis of your parabola has to pass at (-1,3), than the x-coordinate of the vertex has to be equal to -1

OpenStudy (anonymous):

I was just looking at the x and y values

OpenStudy (michele_laino):

 for example or: